# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 4.4 Simple applications to deeply intransitive operator spaces
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$. Let $\mathcal{S}$ be a target-reduced $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Assume that $\operatorname{\operatorname{trk}}\mathcal{S}=nd-1$ and $|\mathbb{F}|\geq nd$. Then $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$.
@@
We choose respective bases $\mathbf{B}_{V}=(e_{1},\dots,e_{nd})$ and $\mathbf{B}_{\mathcal{S}}=(f_{1},\dots,f_{N})$ of $V$ and $\mathcal{S}$ seen as $\mathbb{F}$-vector spaces, and we consider the generic matrix $\mathbf{M}$ associated with $x\in U\mapsto\widehat{x}\in\widehat{\mathcal{S}}$ in these bases. The condition that $\operatorname{\operatorname{trk}}\mathcal{S}=nd-1$ yields $\operatorname{\operatorname{rk}}\mathbf{M}=nd-1$, by Corollary 4.2.
Let $b_{1}$ and $b_{2}$ be nonzero alternators of $\mathcal{S}$. We shall prove that $b_{1}$ and $b_{2}$ are linearly dependent over $\mathbb{F}$.
@@
Now, assume that $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}=1$. Then $\mathbf{B_{1}}=\mathbf{b_{1}}B_{1}$ for some nonzero row $B_{1}\in\operatorname{\operatorname{M}}_{1,nd}(\mathbb{F})\smallsetminus\{0\}$ and some $1$-homogeneous element $\mathbf{b_{1}}$ of $R\smallsetminus\{0\}$. Then the identity $\mathbf{B_{1}}\mathbf{M}=0$ leads to $\mathbf{b_{1}}(B_{1}\mathbf{M})=0$, and hence to $B_{1}\mathbf{M}=0$. The matrix $B_{1}$ represents a linear form $\varphi$ in the dual basis of $\mathbf{B}_{V}$, and the statement that $B_{1}\mathbf{M}=0$ means that all the operators in $\widehat{\mathcal{S}}$ map into the kernel of $\varphi$. Hence so do all the operators in $\mathcal{S}$, and we contradict the assumption that $\mathcal{S}$ is target-reduced.
Hence $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}>1$ and we deduce from the Factorization Lemma that $\mathbf{B_{2}}=\alpha\mathbf{B}_{1}$ for some scalar $\alpha\in\mathbb{F}$. This yields $b_{2}=\alpha b_{1}$. We conclude that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension at most $1$. ∎
@@
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$. Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that $|\mathbb{F}|\geq nd$.
@@
    If $\mathcal{S}$ is weakly primitively intransitive, then it is deeply intransitive.
@@
    If $\mathcal{S}$ is primitively intransitive, then it is deeply intransitive.
@@
Assume that $\mathcal{S}$ is not deeply intransitive. Then we can find a nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that $\mathcal{S}^{V^{\prime}}$ is not intransitive. Denote by $\pi:V\twoheadrightarrow V/V^{\prime}$ the standard projection, and consider the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$.
Take a basis $(v_{1},\dots,v_{s})$ of $\mathcal{S}^{V^{\prime}}$; and lift a basis of $\pi\mathcal{S}$ to a family $(u_{1},\dots,u_{r})$ in $\mathcal{S}$, so that $(u_{1},\dots,u_{r},v_{1},\dots,v_{s})$ is a basis $\mathbf{B}_{\mathcal{S}}$ of $\mathcal{S}$. Take an arbitrary $\mathbb{F}$-basis $\mathbf{B}_{V}$ of $V$ that is adapted to $V^{\prime}$ and in which the first $d$ vectors are pairwise linearly dependent over $\mathbb{D}$, and consider a generic matrix $\mathbf{M}$ of $\widehat{\mathcal{S}}$ in the bases $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$. Then this matrix takes the form
@@
Note that $\mathbf{C}$ is a generic matrix of $\widehat{\mathcal{S}^{V^{\prime}}}$, whereas $\mathbf{B}$ is a generic matrix of $\widehat{\pi\mathcal{S}}$. Hence $\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{trk}}(\mathcal{S}^{V^{\prime}})=\dim_{\mathbb{F}}V^{\prime}=p$. Besides $\operatorname{\operatorname{rk}}\mathbf{C}+\operatorname{\operatorname{rk}}\mathbf{B}\leq\operatorname{\operatorname{rk}}\mathbf{M}$.
Consider the $\mathbb{D}$-linear subspace $L$ of $V$ spanned by the first $d$ vectors of $\mathbf{B}_{V}$, and the matrix $\mathbf{M^{\prime}}$ obtained by deleting the first $d$ rows of $\mathbf{M}$. Then $\mathbf{M^{\prime}}$ is a generic matrix of $\pi^{\prime}\mathcal{S}$ where $\pi^{\prime}:V\twoheadrightarrow V/L$ is the standard projection. By deleting the first $d$ rows we see that
@@
Hence $\operatorname{\operatorname{trk}}(\pi^{\prime}\mathcal{S})\leq\operatorname{\operatorname{trk}}(\mathcal{S})-d$. We conclude that $\mathcal{S}$ is not weakly primitively intransitive.
@@
5 Nondegenerate alternators of large deeply intransitive operator spaces
@@
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim V$. Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
Then $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is $1$-dimensional.
@@ 5.1 Setup
Throughout, we let $n\geq 1$ be an integer, and $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}V=n$. We assume that $|\mathbb{F}|\geq nd$. For simplicity, we set $\varepsilon=1$ if $d=1$, and $\varepsilon=0$ otherwise. Finally, we take a deeply intransitive $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that
$\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon.$
@@
  • In any case $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$.
  • If $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ then $\mathcal{S}$ has a nonzero alternator.
Since $n>0$, we immediately retrieve from Corollary 3.5 that $\mathcal{S}$ is target-reduced.
@@
$r:=\operatorname{\operatorname{trk}}(\mathcal{S})\in\mathopen{[\![}0,nd-1\mathclose{]\!]}.$
Since $\mathcal{S}$ is target-reduced we have $\mathcal{S}\neq\{0\}$ and hence $r>0$. We fix a vector $x\in U\smallsetminus\{0\}$ such that $\operatorname{\operatorname{rk}}\widehat{x}=r$. Then we apply the Tangent Space Lemma to recover that all the operators in
$\mathcal{S}^{\prime}:=\{u\in\mathcal{S}:\;u(x)=0\}$
map into $\mathcal{S}x$. We denote by $H$ the socle of $\mathcal{S}x$ and obtain that all the operators in $\mathcal{S}^{\prime}$ map into $H$ (despite the notation, at this point $H$ is not necessarily a linear hyperplane of the $\mathbb{D}$-vector space $V$). We set
@@
Note that $\mathcal{S}^{\prime}$ is included in $\mathcal{S}^{H}$, which is deeply intransitive, and hence $\mathcal{S}^{\prime}$ is also a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
Now it is time to represent the dual operator space $\widehat{\mathcal{S}}$ in appropriate bases of $\mathcal{S}$ and $V$. We will choose such bases carefully with respect to the vector $x$. For the basis of $V$:
  • We start from a basis $(e_{1},\dots,e_{r})$ of $\mathcal{S}x$ in which the first $s$ vectors span $H$ (as an $\mathbb{F}$-linear subspace), and we extend it arbitrarily into a basis $\mathbf{E}$ of $V$.
  • We lift the first vectors $e_{1},\dots,e_{r}$ to obtain a family $(u_{1},\dots,u_{r})$ of operators in $\mathcal{S}$, so that $u_{i}(x)=e_{i}$ for all $i\in\mathopen{[\![}1,r\mathclose{]\!]}$. Then $(u_{1},\dots,u_{r})$ is linearly independent and we can choose an arbitrary basis $(u_{r+1},\dots,u_{\dim\mathcal{S}})$ of $\mathcal{S}^{\prime}$. Then our basis of $\mathcal{S}$ is $\mathbf{F}:=(u_{1},\dots,u_{r},u_{r+1},\dots,u_{\dim\mathcal{S}})$.
We consider the matrix space $\mathcal{M}$ that represents $\widehat{\mathcal{S}}$ in the bases $\mathbf{F}$ and $\mathbf{E}$. The matrix in $\mathcal{M}$ that corresponds to the operator $\widehat{x}$ is the reduced matrix
$J_{r}=\begin{bmatrix}I_{r}&[0]_{r\times(\dim\mathcal{S}-r)}\\ [0]_{(nd-r)\times r}&[0]_{(nd-r)\times(\dim\mathcal{S}-r)}\end{bmatrix}$.
Now, we take $E:=U$ as our parameter space, $R$ and $\mathbb{L}$ as the associated ring and field, and we construct a generic matrix $\mathbf{M}$ of $\mathcal{M}$ as follows: for all $(i,j)\in\mathopen{[\![}1,nd\mathclose{]\!]}\times\mathopen{[\![}1,\dim\mathcal{S}\mathclose{]\!]}$, we define $\mathbf{m}_{i,j}$ as the linear form on $U$ that takes every vector $z\in U$ to the $(i,j)$-entry of the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$. Hence, the specialization $\mathbf{M}(z)$ is simply the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
As we already know that all the operators in $\mathcal{S}^{\prime}$ map into $H$, the matrix $\mathbf{M}$ actually takes the refined form
$\mathbf{M}=\begin{bmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{B}&[0]_{(nd-r)\times(\dim\mathcal{S}-r)}\end{bmatrix}$
for some matrices $\mathbf{A}\in\operatorname{\operatorname{M}}_{r}(R)$, $\mathbf{B}\in\operatorname{\operatorname{M}}_{nd-r,r}(R)$ and $\mathbf{C}\in\operatorname{\operatorname{M}}_{r,\dim\mathcal{S}-r}(R)$ with $1$-homogeneous entries, and we can further refine
$\mathbf{C}=\begin{bmatrix}\mathbf{C_{1}}\\ [0]_{(r-s)\times(\dim S-r)}\end{bmatrix}\qquad\text{with $\mathbf{C_{1}}\in\operatorname{\operatorname{M}}_{s,\dim\mathcal{S}-r}(R)$.}$
@@
Critically, $\mathbf{C_{1}}$ is a generic matrix for the dual operator space $\widehat{\mathcal{S}^{\prime}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S}^{\prime},H)$ in the basis $(u_{r+1},\dots,u_{\dim\mathcal{S}})$ of $\mathcal{S}^{\prime}$ and the basis $(e_{1},\dots,e_{s})$ of the $\mathbb{F}$-vector space $H$. Here, the parameter space for this generic matrix remains $U$.
@@
Before we proceed, it is time to complete the special case $n=1$ for the proof that $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$. So, assume for a moment that $n=1$. In that case $\mathcal{S}^{\prime}=\{0\}$ and hence $u\in\mathcal{S}\mapsto u(x)\in\mathcal{S}x$ is an $\mathbb{F}$-linear isomorphism. It follows that $r=\dim\mathcal{S}$. The assumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon$ requires here that $\dim\mathcal{S}\geq nd-1$ and hence r=nd1r=nd-1.
From now on, we assume that $n\geq 2$. We will now draw some consequences of the assumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon$.
@@
Assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1. Then the space $H$ has dimension n1n-1 over $\mathbb{D}$, and the space $\mathcal{S}^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$ is target-reduced.
@@
Assume that $\dim_{\mathbb{D}}H<n-1$ or that $\mathcal{S}^{\prime}$ is not target-reduced. Recall from Remark 3.1 that this yields a $\mathbb{D}$-linear subspace $G$ of $V$ with dimension n2n-2 such that $\mathcal{S}^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,G)$. Since $\mathcal{S}^{\prime}$ is deeply intransitive we deduce from point (a) of Theorem 1.20 that $\dim\mathcal{S}^{\prime}\leq\alpha(n-2,d)$. Hence $\dim\mathcal{S}-r\leq\alpha(n-2,d)$. Yet the starting assumption shows that $\dim\mathcal{S}-r\geq\alpha(n-1,d)-n+2+\varepsilon$, and hence $\alpha(n-2,d)\geq\alpha(n-1,d)-n+2+\varepsilon$, to the effect that $(n-1)d-1=\alpha(n-1,d)-\alpha(n-2,d)\leq n-2-\varepsilon$. Hence $(n-1)(d-1)\leq-\varepsilon$, which contradicts the definition of $\varepsilon$ because $n\geq 2$. ∎
@@
The $\mathbb{D}$-linear mappings $u_{1},\dots,u_{(n-1)d}$ map into $H^{\prime}$, and so do all the elements of $\mathcal{S}^{\prime}$. It follows that $\dim(\mathcal{S}^{H^{\prime}})\geq\dim\mathcal{S}^{\prime}+(n-1)d$. Yet $\mathcal{S}^{H^{\prime}}$ is deeply intransitive, so we can apply point (a) of Theorem 1.20 once more to find $\dim\mathcal{S}^{\prime}+(n-1)d\leq\alpha(n-1,d)$, leading to $\dim\mathcal{S}\leq\alpha(n-1,d)-(n-1)d+r=\alpha(n-2,d)+r-1$. A contradiction is then found like in the proof of Claim 5.1, using only $\dim\mathcal{S}-r\geq\alpha(n-1,d)-n+1+\varepsilon$ here. ∎
@@
Assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1. Let $X\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$. Then $\operatorname{\operatorname{sprk}}(X\mathbf{B_{1}})>1$.
@@
This contradicts the fact that $\mathcal{S}^{\prime}$ is target-reduced (see Claim 5.1). ∎

### v2

@@ 4.4 Simple applications to deeply intransitive operator spaces
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$. Let $S$ be a target-reduced $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Assume that $\operatorname{\operatorname{trk}}S=nd-1$ and $|\mathbb{F}|\geq nd$. Then $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$.
@@
We choose respective bases $\mathbf{B}_{V}=(e_{1},\dots,e_{nd})$ and $\mathbf{B}_{S}=(f_{1},\dots,f_{N})$ of $V$ and $S$ seen as $\mathbb{F}$-vector spaces, and we consider the generic matrix $\mathbf{M}$ associated with $x\in U\mapsto\widehat{x}\in\widehat{S}$ in these bases. The condition that $\operatorname{\operatorname{trk}}S=nd-1$ yields $\operatorname{\operatorname{rk}}\mathbf{M}=nd-1$, by Corollary 4.2.
Let $b_{1}$ and $b_{2}$ be nonzero alternators of $S$. We shall prove that $b_{1}$ and $b_{2}$ are linearly dependent over $\mathbb{F}$.
@@
Now, assume that $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}=1$. Then $\mathbf{B_{1}}=\mathbf{b_{1}}B_{1}$ for some nonzero row $B_{1}\in\operatorname{\operatorname{M}}_{1,nd}(\mathbb{F})\smallsetminus\{0\}$ and some $1$-homogeneous element $\mathbf{b_{1}}$ of $R\smallsetminus\{0\}$. Then the identity $\mathbf{B_{1}}\mathbf{M}=0$ leads to $\mathbf{b_{1}}(B_{1}\mathbf{M})=0$, and hence to $B_{1}\mathbf{M}=0$. The matrix $B_{1}$ represents a linear form $\varphi$ in the dual basis of $\mathbf{B}_{V}$, and the statement that $B_{1}\mathbf{M}=0$ means that all the operators in $\widehat{S}$ map into the kernel of $\varphi$. Hence so do all the operators in $S$, and we contradict the assumption that $S$ is target-reduced.
Hence $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}>1$ and we deduce from the Factorization Lemma that $\mathbf{B_{2}}=\alpha\mathbf{B}_{1}$ for some scalar $\alpha\in\mathbb{F}$. This yields $b_{2}=\alpha b_{1}$. We conclude that $\operatorname{\operatorname{Alt}}(S)$ has dimension at most $1$. ∎
@@
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$. Let $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that $|\mathbb{F}|\geq nd$.
@@
    If $S$ is weakly primitively intransitive, then it is deeply intransitive.
@@
    If $S$ is primitively intransitive, then it is deeply intransitive.
@@
Assume that $S$ is not deeply intransitive. Then we can find a nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that $S^{V^{\prime}}$ is not intransitive. Denote by $\pi:V\twoheadrightarrow V/V^{\prime}$ the standard projection, and consider the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$.
Take a basis $(v_{1},\dots,v_{s})$ of $S^{V^{\prime}}$; and lift a basis of $\pi S$ to a family $(u_{1},\dots,u_{r})$ in $S$, so that $(u_{1},\dots,u_{r},v_{1},\dots,v_{s})$ is a basis $\mathbf{B}_{S}$ of $S$. Take an arbitrary $\mathbb{F}$-basis $\mathbf{B}_{V}$ of $V$ that is adapted to $V^{\prime}$ and in which the first $d$ vectors are pairwise linearly dependent over $\mathbb{D}$, and consider a generic matrix $\mathbf{M}$ of $\widehat{S}$ in the bases $\mathbf{B}_{S}$ and $\mathbf{B}_{V}$. Then this matrix takes the form
@@
Note that $\mathbf{C}$ is a generic matrix of $\widehat{S^{V^{\prime}}}$, whereas $\mathbf{B}$ is a generic matrix of $\widehat{\pi S}$. Hence $\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{trk}}(S^{V^{\prime}})=\dim_{\mathbb{F}}V^{\prime}=p$. Besides $\operatorname{\operatorname{rk}}\mathbf{C}+\operatorname{\operatorname{rk}}\mathbf{B}\leq\operatorname{\operatorname{rk}}\mathbf{M}$.
Consider the $\mathbb{D}$-linear subspace $L$ of $V$ spanned by the first $d$ vectors of $\mathbf{B}_{V}$, and the matrix $\mathbf{M^{\prime}}$ obtained by deleting the first $d$ rows of $\mathbf{M}$. Then $\mathbf{M^{\prime}}$ is a generic matrix of $\pi^{\prime}S$ where $\pi^{\prime}:V\twoheadrightarrow V/L$ is the standard projection. By deleting the first $d$ rows we see that
@@
Hence $\operatorname{\operatorname{trk}}(\pi^{\prime}S)\leq\operatorname{\operatorname{trk}}(S)-d$. We conclude that $S$ is not weakly primitively intransitive.
@@
5 Nondegenerate alternators of deeply intransitive operator spaces
@@
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim V$. Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$.
Then $\operatorname{\operatorname{Alt}}(S)$ is $1$-dimensional.
@@ 5.1 Setup
Throughout, we let $n\geq 1$ be an integer, and $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}V=n$. We assume that $|\mathbb{F}|\geq nd$. For simplicity, we set $\varepsilon=1$ if $d=1$, and $\varepsilon=0$ otherwise. Finally, we take a deeply intransitive $\mathbb{F}$-linear subspace $S$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that
$\dim S\geq\alpha(n,d)-n+1+\varepsilon.$
@@
  • In any case $\operatorname{\operatorname{trk}}(S)=nd-1$.
  • If $\dim S\geq\alpha(n,d)-n+2+\varepsilon$ then $S$ has a nonzero alternator.
Since $n>0$, we immediately retrieve from Corollary 3.5 that $S$ is target-reduced.
@@
$r:=\operatorname{\operatorname{trk}}(S)\in\mathopen{[\![}0,nd-1\mathclose{]\!]}.$
Since $S$ is target-reduced we have $S\neq\{0\}$ and hence $r>0$. We fix a vector $x\in U\smallsetminus\{0\}$ such that $\operatorname{\operatorname{rk}}\widehat{x}=r$. Then we apply the Tangent Space Lemma to recover that all the operators in
$S^{\prime}:=\{u\in S:\;u(x)=0\}$
map into $Sx$. We denote by $H$ the socle of $Sx$ and obtain that all the operators in $S^{\prime}$ map into $H$ (despite the notation, at this point $H$ is not necessarily a linear hyperplane of the $\mathbb{D}$-vector space $V$). We set
@@
Note that $S^{\prime}$ is included in $S^{H}$, which is deeply intransitive, and hence $S^{\prime}$ is also a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
Now it is time to represent the dual operator space $\widehat{S}$ in appropriate bases of $S$ and $V$. We will choose such bases carefully with respect to the vector $x$. For the basis of $V$:
  • We start from a basis $(e_{1},\dots,e_{r})$ of $Sx$ in which the first $s$ vectors span $H$ (as an $\mathbb{F}$-linear subspace), and we extend it arbitrarily into a basis $\mathbf{E}$ of $V$.
  • We lift the first vectors $e_{1},\dots,e_{r}$ to obtain a family $(u_{1},\dots,u_{r})$ of operators in $S$, so that $u_{i}(x)=e_{i}$ for all $i\in\mathopen{[\![}1,r\mathclose{]\!]}$. Then $(u_{1},\dots,u_{r})$ is linearly independent and we can choose an arbitrary basis $(u_{r+1},\dots,u_{\dim S})$ of $S^{\prime}$. Then our basis of $S$ is $\mathbf{F}:=(u_{1},\dots,u_{r},u_{r+1},\dots,u_{\dim S})$.
We consider the matrix space $\mathcal{M}$ that represents $\widehat{S}$ in the bases $\mathbf{F}$ and $\mathbf{E}$. The matrix in $\mathcal{M}$ that corresponds to the operator $\widehat{x}$ is the reduced matrix
$J_{r}=\begin{bmatrix}I_{r}&[0]_{r\times(\dim S-r)}\\ [0]_{(nd-r)\times r}&[0]_{(nd-r)\times(\dim S-r)}\end{bmatrix}$.
Now, we take $E:=U$ as our parameter space, $R$ and $\mathbb{L}$ as the associated ring and field, and we construct a generic matrix $\mathbf{M}$ of $\mathcal{M}$ as follows: for all $(i,j)\in\mathopen{[\![}1,nd\mathclose{]\!]}\times\mathopen{[\![}1,\dim S\mathclose{]\!]}$, we define $\mathbf{m}_{i,j}$ as the linear form on $U$ that takes every vector $z\in U$ to the $(i,j)$-entry of the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$. Hence, the specialization $\mathbf{M}(z)$ is simply the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
As we already know that all the operators in $S^{\prime}$ map into $H$, the matrix $\mathbf{M}$ actually takes the refined form
$\mathbf{M}=\begin{bmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{B}&[0]_{(nd-r)\times(\dim S-r)}\end{bmatrix}$
for some matrices $\mathbf{A}\in\operatorname{\operatorname{M}}_{r}(R)$, $\mathbf{B}\in\operatorname{\operatorname{M}}_{nd-r,r}(R)$ and $\mathbf{C}\in\operatorname{\operatorname{M}}_{r,\dim S-r}(R)$ with $1$-homogeneous entries, and we can further refine
$\mathbf{C}=\begin{bmatrix}\mathbf{C_{1}}\\ [0]_{(r-s)\times(\dim S-r)}\end{bmatrix}\qquad\text{with $\mathbf{C_{1}}\in\operatorname{\operatorname{M}}_{s,\dim S-r}(R)$.}$
@@
Critically, $\mathbf{C_{1}}$ is a generic matrix for the dual operator space $\widehat{S^{\prime}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S^{\prime},H)$ in the basis $(u_{r+1},\dots,u_{\dim S})$ of $S^{\prime}$ and the basis $(e_{1},\dots,e_{s})$ of the $\mathbb{F}$-vector space $H$. Here, the parameter space for this generic matrix remains $U$.
@@
Before we proceed, it is time to complete the special case $n=1$ for the proof that $\operatorname{\operatorname{trk}}(S)=nd-1$. So, assume for a moment that $n=1$. In that case $S^{\prime}=\{0\}$ and hence $u\in S\mapsto u(x)\in Sx$ is an $\mathbb{F}$-linear isomorphism. It follows that $r=\dim S$. The assumption that $\dim S\geq\alpha(n,d)-n+1+\varepsilon$ requires here that $\dim S\geq nd-1$ and hence r=nd1r=nd-1.
From now on, we assume that $n\geq 2$. We will now draw some consequences of the assumption that $\dim S\geq\alpha(n,d)-n+1+\varepsilon$.
@@
Assume that $\dim S\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1. Then the space $H$ has dimension n1n-1 over $\mathbb{D}$, and the space $S^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$ is target-reduced.
@@
Assume that $\dim_{\mathbb{D}}H<n-1$ or that $S^{\prime}$ is not target-reduced. Recall from Remark 3.1 that this yields a $\mathbb{D}$-linear subspace $G$ of $V$ with dimension n2n-2 such that $S^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,G)$. Since $S^{\prime}$ is deeply intransitive we deduce from point (a) of Theorem 1.22 that $\dim S^{\prime}\leq\alpha(n-2,d)$. Hence $\dim S-r\leq\alpha(n-2,d)$. Yet the starting assumption shows that $\dim S-r\geq\alpha(n-1,d)-n+2+\varepsilon$, and hence $\alpha(n-2,d)\geq\alpha(n-1,d)-n+2+\varepsilon$, to the effect that $(n-1)d-1=\alpha(n-1,d)-\alpha(n-2,d)\leq n-2-\varepsilon$. Hence $(n-1)(d-1)\leq-\varepsilon$, which contradicts the definition of $\varepsilon$ because $n\geq 2$. ∎
@@
The $\mathbb{D}$-linear mappings $u_{1},\dots,u_{(n-1)d}$ map into $H^{\prime}$, and so do all the elements of $S^{\prime}$. It follows that $\dim(S^{H^{\prime}})\geq\dim S^{\prime}+(n-1)d$. Yet $S^{H^{\prime}}$ is deeply intransitive, so we can apply point (a) of Theorem 1.22 once more to find $\dim S^{\prime}+(n-1)d\leq\alpha(n-1,d)$, leading to $\dim S\leq\alpha(n-1,d)-(n-1)d+r=\alpha(n-2,d)+r-1$. A contradiction is then found like in the proof of Claim 5.1, using only $\dim S-r\geq\alpha(n-1,d)-n+1+\varepsilon$ here. ∎
@@
Assume that $\dim S\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1. Let $X\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$. Then $\operatorname{\operatorname{sprk}}(X\mathbf{B_{1}})>1$.
@@
This contradicts the fact that $S^{\prime}$ is target-reduced (see Claim 5.1). ∎